what are normal ranges of temp in centigrade and fahrenheit, blood pressure, respiration, pulse rates and oxygen saturation for all major age groups?
Answer #1

Normal Values By Age Age Temp Pulse Average Respirations Blood Degrees And Ranges Averages and Averages and Pressure Celsius

Formula Example Convert Celcius to Fahrenheit F= 1.8C + 32 C= 37 F= (1.8)(37) + 32 F= 66.6 + 32 F= 98.6°

Normal arterial blood oxygen saturation levels in humans are 95–100 percent. If the level is below 90 percent, it is considered low resulting in hypoxemia.

KINDLEY,LIKE&SUPPORT